OpenStudy (anonymous):

What are the solutions to P(x) = x4 + 2x3 + 6x2 - 3x - 7

OpenStudy (anonymous):

use substitution let y = x^2 now you would have p(y) = y^2 + 2y + 3-7 now combine the like terms and factor, or use quadratic equation

OpenStudy (dumbcow):

the only possible rational solutions are +-1, +-7 if one of these works you can factor it out resulting in a cubic function....repeat until you get a quadratic graphing it on a graphing calculator is also helpful :)

OpenStudy (anonymous):

@dumbcow I tested the factors +-1 and +-7 using synthetic division but none of them work.

OpenStudy (dumbcow):

that means solutions are irrational, unless you can factor the polynomial then only way to get solutions is by approximating using calculator or more complex methods involving calculus

OpenStudy (anonymous):

So will the solutions dhow up on a graph?

OpenStudy (anonymous):

Because I have to describe a graph mainly focusing on the end behaviors and the x-intercepts but I can't find the solutions to the equation